You wake up from a nightmare in which you've completely forgotten about the final exam in your stats class. Since you haven't studied, you estimate that there is an 80% chance of answering any individual problem correctly. The test is 5 questions long. If you wanted to find the probability you answer all the questions incorrectly, what could you do?
A) 0.8 × 0.8 × 0.8 ×0.8 × 0.8
B) 0.2 + 0.2+0.2+0.2+0.2
C) 1 - 0.8
D) 0.2×0.2×0.2×0.2×0.2
Answer #1

as answering incorrectly a problem =1-0.8 =0.2

therefore due to independence: P(answer all the questions incorrectly) =0.2*0.2*0.2*0.2*0.2

option D is correct'